OpenStudy (anonymous):

(2^m) (5^n) = 400,000 m = ? n = ?

OpenStudy (anonymous):

I don't know how to start solving this, could someone show me how?

OpenStudy (anonymous):

...Help?

OpenStudy (anonymous):

\[400,000=4*10^5=2^2*(2*5)^5=2^2*2^5*5^5=?\]

OpenStudy (anonymous):

oh I see